# One Model for All Tasks: Leveraging Efficient World Models in Multi-Task Planning

### v1

@@ Abstract
In heterogeneous multi-task learning, tasks not only exhibit diverse observation and action spaces but also vary substantially in intrinsic difficulty. While conventional multi-task world models like UniZero excel in single-task settings, we find that when handling large-scale heterogeneous environments, gradient conflicts and the loss of model plasticity often constrain their sample and computational efficiency. In this work, we address these challenges from two perspectives: the single learning iteration and the overall learning process. First, we investigate the impact of key design spaces on extending UniZero to multi-task planning. We find that a Mixture-of-Experts (MoE) architecture provides the most substantial performance gains by mitigating gradient conflicts, leading to our proposed model, ScaleZero. Second, to dynamically balance the computational load across the learning process, we introduce an online, LoRA-based dynamic parameter scaling (DPS) strategy. This strategy progressively integrates LoRA adapters in response to task-specific progress, enabling adaptive knowledge retention and parameter expansion. Empirical evaluations on standard benchmarks such as Atari, DMControl (DMC), and Jericho demonstrate that ScaleZero, relying exclusively on online reinforcement learning with one model, attains performance on par with specialized single-task baselines. Furthermore, when augmented with our dynamic parameter scaling strategy, our method achieves competitive performance while requiring only 80% of the single-task environment interaction steps. These findings underscore the potential of ScaleZero for effective large-scale multi-task learning. Our code is available at https://github.com/opendilab/LightZero.
@@ 1 Introduction
Unified world models (Reed et al., 2022; Gallouédec et al., 2024b; Lee et al., 2022b) represent a significant step towards generalist agents, offering a single, cohesive framework for multi-modal perception, long-horizon prediction, and decision-making by learning a shared latent space representation. The true power of these models is unlocked when combined with search-based planning algorithms like Monte Carlo Tree Search (MCTS) (Silver et al., 2017; Schrittwieser et al., 2019), which can perform efficient lookahead within this compact latent space. MCTS excels at dynamically balancing exploration and exploitation in complex decision spaces, a combination that has achieved well-established success in homogeneous task domains like board games (Ye et al., 2021; Silver et al., 2017; Schrittwieser et al., 2019). The natural next frontier is to extend this success to heterogeneous multi-task reinforcement learning (MTRL) (Yu et al., 2024), aiming to build a single, generalist agent that can master a wide array of diverse environments. However, this ambition is impeded by a formidable obstacle: training a single shared model on diverse tasks with potentially conflicting dynamics and objectives is notoriously difficult. While prior work (Fernando et al., 2023; Shi et al., 2023) has attempted to mitigate inter-task interference, these methods have been predominantly studied in supervised learning settings. The challenges and dynamics of multi-task planning within a shared world model remain largely unexplored.
@@
Our investigation into this specific domain of multi-task planning identifies and addresses two critical, intertwined obstacles that are not fully resolved by existing approaches: (1) Representational bottlenecks and plasticity collapse: In diverse multi-task settings, a shared model is susceptible to gradient dominance from simpler, faster-converging tasks (Cho et al., 2024). This imbalance leads to representation interference (Yu et al., 2020; Bejnordi et al., 2024), where the learning signals for more complex tasks are suppressed. The consequence is a progressive loss of network plasticity (Dohare et al., 2024; Todorov et al., 2025)the fundamental ability of a model to adjust its parameters to learn from new data–imposing a hard ceiling on the model’s overall learning capacity and leading to performance collapse on challenging tasks. (2) Static resource allocation: Conventional architectures employ a uniform, one-size-fits-all forward pass, applying the same computational effort to every task irrespective of its intrinsic difficulty. This static strategy results in profound computational inefficiency, as resources are squandered on already-mastered tasks instead of being directed toward those that require further learning.
To dissect these issues, we first conduct a rigorous empirical diagnosis in Figure 1. By deploying a baseline UniZero model across eight canonical Atari games, we observe a stark failure pattern. While the model rapidly masters simple tasks like Pong, it exhibits initial progress followed by a catastrophic performance collapse on complex, visually distinct games like Seaquest and ChopperCommand. We provide the quantitative evidence linking this failure to specific internal model dynamics: an uncontrolled inflation of the latent state norm and a corresponding spike in the dormant neuron ratio within the Transformer backbone, signaling a descent into a "frozen," non-adaptive state.
@@
Motivated by this diagnosis, we undertake a principled exploration of the UniZero design space across the axes of input, model, and optimization from the the single learning iteration perspective. This systematic investigation, which evaluates task embeddings, encoder architectures (ResNet (He et al., 2016) vs. ViT (Dosovitskiy et al., 2020)), latent normalization schemes (Hansen et al., 2023), sparse Mixture-of-Experts (MoE) backbones, and gradient rectification methods(He et al., 2020), culminates in a new, powerful model we term ScaleZero. Validated on a comprehensive suite of online RL benchmarks—spanning 26 Atari games, 18 DeepMind Control tasks, and 4 Jericho environments with different observation and action space—ScaleZero consistently matches and often surpasses the performance of specific single-task expert agents.
@@
Finally, to resolve the challenge of computational inefficiency, we introduce Dynamic Parameter Scaling (DPS) from the overall learning process perspective, a novel online mechanism that couples model capacity to learning progress. DPS adaptively curates the set of active tasks based on real-time return feedback and orchestrates a phased expansion of model capacity by injecting lightweight LoRA adapters. By strategically freezing previously trained parameters, DPS creates a curriculum of model complexity that directs computational resources where they are most needed. Our experiments demonstrate that when augmented with our dynamic parameter scaling strategy, our method achieves performance nearly on par with single-task agents while reducing total environment interactions by around 20%, offering a superior trade-off between final performance and computational budget.
@@
Our main contributions are summarized as follows: (1) We provide the quantitative diagnosis of plasticity collapse in unified world models within heterogeneous MTRL, establishing a concrete link between performance degradation and internal network dynamics. (2) We conduct a systematic architectural exploration that yields ScaleZero, a unified world model that demonstrates exceptional performance and generalization across distinct tasks from Atari (diverse visual challenges), DMControl (state-based continuous control), and Jericho (text-based compositional reasoning) benchmark. (3) We propose Dynamic Parameter Scaling (DPS), an adaptive training strategy that dynamically allocates model capacity and computational resources, reducing total environment interactions by around 20%.
@@ 2 Background
2.1 Unified World Models in MCTS-Based Reinforcement Learning
@@
Model-based reinforcement learning, particularly methods employing MCTS for planning within a learned latent space, has set the standard in complex sequential decision-making domains (Schrittwieser et al., 2019; Antonoglou et al., 2021; Ye et al., 2021). These methods learn a world model composed of three core components: (i) a representation model, $z_{t}=h_{\theta}(o_{t-H+1:t},a_{t-H+1:t-1})$, which encodes an observation history into a latent state, where $H$ is the training sequence length; (ii) a dynamics model, $(\hat{z}_{t+1},\hat{r}_{t})=g_{\theta}(z_{t},a_{t})$, predicting the subsequent latent state and reward; and (iii) a prediction model, $(p_{t},v_{t})=f_{\theta}(z_{t})$, which estimates a policy and value to guide the MCTS planner. Recent architectures like UniZero (Pu et al., 2024) push this paradigm further by unifying these three components into a single Transformer-based architecture. This monolithic design allows for end-to-end optimization via a composite loss function that aligns the model’s predictions with MCTS-derived targets for policy ($\pi_{t}$) and value ($\hat{v}_{t}$), alongside objectives for reward and state reconstruction:
@@
$\mathcal{L}_{\text{UniZero}}=\sum_{t=0}^{H-1}\left(\mathcal{L}_{\text{value}}(v_{t},\hat{v}_{t})+\mathcal{L}_{\text{policy}}(p_{t},\pi_{t})+\mathcal{L}_{\text{reward}}(\hat{r}_{t},r_{t})+\mathcal{L}_{\text{dynamics}}(\hat{z}_{t+1},\operatorname{sg}(z_{t+1}))\right),$ (1)
where $\operatorname{sg}(\cdot)$ denotes the stop-gradient operator. While parameter-efficient, this unified training scheme creates a critical vulnerability in heterogeneous settings: the shared parameters are subject to a single, aggregated gradient signal, which can be dominated by easier tasks (Ma et al., 2023), thereby impeding learning on more complex ones and leading to plasticity collapse.
2.2 Gradient Conflict in Multi-Task Reinforcement Learning
@@
In MTRL, we consider a distribution of $K$ distinct tasks, $\mathcal{T}=\{\tau_{1},\dots,\tau_{K}\}$. The objective is to learn a single set of shared parameters $\theta$ for a task-conditioned policy $\pi_{\theta}(a|s,k)$ that maximizes expected returns across all tasks, where $k\in\{1,\dots,K\}$ is the task identifier. Given a batch of data from all tasks, the total loss is the sum of the individual task losses: $\mathcal{L}_{\text{MTRL}}(\theta)=\sum_{k=1}^{K}\mathcal{L}^{k}_{\text{UniZero}},$ where $\mathcal{L}^{k}_{\text{UniZero}}$ is the loss computed on data from task $k$. The learning dynamics are driven by the total gradient $G_{\text{total}}=\nabla_{\theta}\mathcal{L}_{\text{MTRL}}$, which is the sum of per-task gradients: $G_{\text{total}}=\sum_{k=1}^{K}g_{k}$, where $g_{k}=\nabla_{\theta}\mathcal{L}^{k}_{\text{UniZero}}$. The core problem, known as gradient conflict, arises when these per-task gradients are misaligned. For any two tasks $i$ and $j$, the conflict can be quantified by the cosine similarity of their gradients: $\cos(g_{i},g_{j})=\frac{g_{i}\cdot g_{j}}{\|g_{i}\|\|g_{j}\|}$. A negative value indicates that an update improving performance on task $i$ will degrade performance on task $j$. In a monolithic architecture where all parameters $\theta$ are shared, frequent negative similarities lead to destructive interference (Sodhani et al., 2021). This forces the model into a suboptimal compromise, creating a representational bottleneck and impeding mastery of tasks with conflicting or complex objectives (Hansen et al., 2023).
2.3 Sparse and Parameter-Efficient Architectures
A significant challenge in developing unified models for heterogeneous environments is overcoming issues such as representational bottlenecks, plasticity collapse, and the rigidity imposed by static resource allocation. To address these challenges, the literature has explored several sparse and parameter-efficient architectures that operate at different perspectives. One prominent approach involves replacing the standard feed-forward network in a Transformer block (Vaswani et al., 2017) with a sparse Mixture-of-Experts architecture (Shazeer et al., 2017; Liu et al., 2024a). This model comprises $M$ expert networks and a trainable sparse gating function, $G(x)$, which selects a small top-k subset of experts for each input. The forward pass is defined as: $\mathrm{MoE}(x)=\sum_{i=1}^{M}G_{i}(x)\cdot\mathrm{Expert}_{i}(x).$ This conditional computation creates distinct representational pathways for different inputs and significantly improve model performance and efficiency with fewer resources. For parameter-efficient adaptation of large models, Low-Rank Adaptation (LoRA) (Hu et al., 2022) has become a widely adopted technique. LoRA adapts a pre-trained weight matrix $W_{0}\in\mathbb{R}^{d\times m}$ by learning a low-rank update, $\Delta W=BA$, where $B\in\mathbb{R}^{d\times r}$ and $A\in\mathbb{R}^{r\times m}$, with the rank $r\ll\min(d,m)$. The effective weight matrix then operates on an input $x$ as: $W(x)=\bigl{(}W_{0}+\alpha\,BA\bigr{)}x,$ where a learnable scalar $\alpha$ controls the magnitude of the update. By freezing the original weights $W_{0}$ and training only the compact low-rank factors $(A,B)$, LoRA facilitates efficient adaptation and isolates task-specific modifications. In this context, our work fuses these ideas from two perspectives. At the single learning iteration level, a gating function sparsely selects LoRA modules, treating them as experts. At the overall learning process level, we incrementally add new LoRA modules for new tasks—freezing prior ones—to efficiently scale model capacity.
@@ 3 Method
This section systematically explores the core challenges that unified world models face in heterogeneous multi-task learning. We begin in Section 3.1 by quantitatively diagnosing the phenomenon of plasticity collapse in shared-parameter models under a multi-task paradigm, uncovering its underlying mechanisms. Subsequently, in Section 3.2, we systematically dissect the design space of the UniZero model. Through a series of ablation studies, we converge on a new architecture, termed ScaleZero, which effectively mitigates plasticity degradation. Finally, to address the inefficiency of static resource allocation, we introduce Dynamic Parameter Scaling (DPS) in Section 3.3. This strategy adaptively expands the model’s capacity in response to task-specific learning progress, thereby achieving a superior trade-off between computational efficiency and final performance.
3.1 Diagnosing Plasticity Collapse in Multi-Task Learning

### v2

@@ Abstract
In heterogeneous multitask decision-making, tasks not only exhibit diverse observation and action spaces but also vary substantially in their underlying complexities. While conventional multitask world models like UniZero excel in single-task settings, we find that when handling a broad and diverse suite of tasks, gradient conflicts and the loss of model plasticity often constrain their sample efficiency. In this work, we address these challenges from two complementary perspectives: the single learning iteration and the overall learning process. First, to mitigate the gradient conflicts, we systematically investigate key architectural designs for extending UniZero. Our investigation identifies a Mixture-of-Experts (MoE) architecture as the most effective approach. We demonstrate, both theoretically and empirically, that this architecture alleviates gradient conflicts by routing task-specific representations to specialized sub-networks. This finding leads to our proposed model, ScaleZero. Second, to dynamically allocate model capacity throughout the learning process, we introduce an online Dynamic Parameter Scaling (DPS) strategy. This strategy progressively integrates LoRA adapters in response to task-specific progress, enabling adaptive knowledge retention and parameter expansion. Evaluations on a diverse set of standard benchmarks (Atari, DMC, Jericho) demonstrate that ScaleZero, utilizing solely online reinforcement learning with one model, performs on par with specialized single-task agents. With the DPS strategy, it remains competitive while using just 71.5% of the environment interactions. These findings underscore the potential of ScaleZero for effective multitask planning.
@@ 1 Introduction
Unified world models (Reed et al., 2022; Lee et al., 2022; Gallouédec et al., 2024) represent a significant step towards generalist agents, offering a cohesive framework for multi-modal perception  (Zhang et al., 2022), long-horizon prediction (Ni et al., 2024), and decision-making by learning a shared latent space representation. By encoding the environment into a compact latent space, these learned representations enable effective planning. Algorithms like Monte Carlo Tree Search (MCTS) (Świechowski et al., 2023) are well-suited for this task, as they can perform lookahead efficiently within this abstract representation (Schrittwieser et al., 2019; 2021). MCTS excels at dynamically balancing exploration and exploitation, a combination that has achieved well-established success in homogeneous task domains like board games (Schrittwieser et al., 2019; Ye et al., 2021). Extending this paradigm to heterogeneous multitask reinforcement learning (MTRL) (Yu et al., 2024) is a key objective for creating generalist agents. However, this ambition is impeded by a formidable obstacle: training a single shared model on diverse tasks with potentially conflicting dynamics and objectives is notoriously difficult (Cho et al., 2024). While prior work (Sun et al., 2022; Fernando et al., 2023; Shi et al., 2023) has attempted to mitigate inter-task interference, these methods have been predominantly studied in supervised learning and model-free RL settings. The challenges and dynamics of multitask planning within a shared world model remain largely unexplored.
@@
Within this specific domain of multitask planning, our work identifies and addresses two critical, intertwined obstacles not fully resolved by existing methods: (1) Representational bottlenecks and plasticity collapse: In diverse multitask settings, a shared model is susceptible to gradient dominance from simpler, faster-converging tasks (Cho et al., 2024). This imbalance leads to representation interference (Yu et al., 2020; Bejnordi et al., 2024), where the learning signals for more complex tasks are suppressed. This culminates in a progressive loss of network plasticity (Dohare et al., 2024; Todorov et al., 2025)the fundamental ability of a model to adjust its parameters to learn from new data—imposing a hard ceiling on the model’s overall learning capacity and leading to performance collapse on challenging tasks. (2) Static resource allocation: Conventional architectures employ a uniform, one-size-fits-all forward pass, applying the same learning resources (e.g., data collection and model updates) to every task irrespective of its intrinsic difficulty. This static strategy results in profound computational inefficiency, as resources are squandered on converged or show diminishing returns, instead of being directed toward those that require further learning. To dissect these issues, we begin with a quantitative diagnosis using UniZero (Pu et al., 2024), a contemporary unified world model, as a representative testbed. We train a single model via online reinforcement learning across eight canonical Atari games and observe a stark failure pattern, as shown in Figure 1. While the model rapidly masters simple tasks like Pong, it exhibits initial progress followed by a catastrophic performance collapse on complex games with disparate dynamics and visual styles, such as Seaquest, whose complex exploration demands starkly contrast with the reactive gameplay of tasks like Pong. We provide quantitative evidence linking this failure to specific internal model dynamics (Lyle et al., 2022): an uncontrolled inflation of the latent state norm and a corresponding spike in the dormant neuron ratio (Sokar et al., 2023) within the transformer backbone, signaling a collapse in model plasticity, rendering the network unable to adapt to new information from challenging tasks.
@@
Based on this diagnosis, we tackle these challenges from two complementary perspectives. First, as an internal, architectural solution, we undertake a principled exploration of the design space across the core dimensions input representation, model architecture, and optimization strategy. This systematic investigation, which evaluates five key axes—task conditioning, encoder architectures (ResNet (He et al., 2016) vs. ViT (Dosovitskiy et al., 2020)), latent normalization schemes, backbone design, and multitask optimization strategies (He et al., 2020)—culminates in a new, powerful model we term ScaleZero. Among the evaluated design choices, the integration of a sparse Mixture-of-Experts (Shazeer et al., 2017) backbone yields the most significant gains by fundamentally addressing plasticity collapse. To provide a deeper understanding of this key architectural choice, we present a dedicated empirical and theoretical analysis in Section 4.3, explaining MoE’s effectiveness in multitask planning. Validated on a comprehensive suite of online RL benchmarks spanning diverse modalities—including 26 visually-complex Atari games (Bellemare et al., 2013), 18 state-based DeepMind Control (DMC) Suite tasks (Tunyasuvunakool et al., 2020), and 4 text-based Jericho environments (Hausknecht et al., 2020)ScaleZero robustly matches and often surpasses the performance of single-task expert agents.
@@
Second, as an external, procedural strategy to optimize the overall learning process, we introduce Dynamic Parameter Scaling (DPS), a novel online mechanism that couples model capacity to learning progress. DPS adaptively curates the set of active tasks based on real-time return feedback and orchestrates a phased expansion of model capacity by injecting lightweight LoRA adapters (Hu et al., 2022). By strategically freezing previously trained parameters, DPS creates a curriculum of model capacity that directs computational resources where they are most needed. Our experiments demonstrate that when augmented with this strategy, our method achieves performance nearly on par with single-task agents while reducing total environment interactions by around 28.5% on the DMC benchmark, offering a superior trade-off between final performance and computational budget.
@@
Our main contributions are summarized as follows: (1) We provide the quantitative diagnosis of plasticity collapse in unified world models within heterogeneous MTRL, establishing a concrete link between performance degradation and internal learning dynamics. (2) We conduct a systematic architectural exploration that yields ScaleZero, a unified world model that demonstrates exceptional performance and generalization across distinct tasks from three distinct benchmarks. (3) We propose Dynamic Parameter Scaling, an adaptive training strategy that dynamically allocates model capacity and computational resources, reducing total environment interactions by around 28.5%.
@@ 2 Background
2.1 MCTS with Learned World Models
@@
Monte Carlo Tree Search is a powerful planning algorithm that has demonstrated remarkable success in domains with known rules (e.g. board games). Methods employing MCTS for planning within a learned latent space represent the state-of-the-art in complex sequential decision-making (Schrittwieser et al., 2019; Ye et al., 2021). These approaches learn a world model comprising three components: (i) a representation model $h_{\theta}$ that encodes observation-action history into a latent state $z_{t}$; (ii) a dynamics model $g_{\theta}$ that predicts the next latent state and reward $(\hat{z}_{t+1},\hat{r}_{t})=g_{\theta}(z_{t},a_{t})$; and (iii) a prediction model $f_{\theta}$ that outputs a policy and value $(p_{t},v_{t})$ to guide the MCTS planner. Recent architectures, exemplified by UniZero (Pu et al., 2024), unify these components into a single, monolithic Transformer. This design enables end-to-end optimization via a composite loss that aligns the model’s predictions with targets derived from the MCTS planner: a policy target ($\pi_{t}$) based on the root node’s visit counts, and a bootstrapped TD value target ($\hat{v}^{targ}_{t}$)  (Schrittwieser et al., 2019). These are combined with objectives for predicting the reward ($r_{t}$) and the next latent state ($z_{t+1}$):
@@
$\mathcal{L}_{\text{Unified}}=\sum_{t=0}^{H-1}\left(\mathcal{L}_{\text{value}}(v_{t},\hat{v}^{targ}_{t})+\mathcal{L}_{\text{policy}}(p_{t},\pi_{t})+\mathcal{L}_{\text{reward}}(\hat{r}_{t},r_{t})+\mathcal{L}_{\text{dynamics}}(\hat{z}_{t+1},\operatorname{sg}(z_{t+1}))\right),$ (1)
where $H$ is the context length, $\operatorname{sg}(\cdot)$ means stop-gradient. While sample-efficient and hyperparameter-insensitive for single-task training, this unified scheme exposes a critical vulnerability: in multitask settings, the shared transformer backbone is updated by an aggregated gradient. This shared update mechanism becomes the nexus of interference, impeding a single model’s ability to learn multiple complex task and leading directly to the plasticity collapse we identified in our diagnosis.
2.2 Gradient Conflict in Unified Architectures
@@
The challenge of MTRL is to train a single, task-conditioned policy $\pi_{\theta}(a|s,k)$ on a distribution of $K$ tasks, $\{\tau_{1},\dots,\tau_{K}\}$. Even within a single task, the composite losses of a world model (e.g., for dynamics, reward, and policy) can conflict, creating a challenging multi-objective optimization problem (Ma et al., 2023). In MTRL, this challenge is compounded by inter-task gradient conflicts. In a unified architecture with shared parameters $\theta$, the total loss is the sum of individual task losses, $\mathcal{L}_{\text{MTRL}}(\theta)=\sum_{k=1}^{K}\mathcal{L}^{k}$, where $\mathcal{L}^{k}$ is the loss from Eq. 1 for task $k$. Consequently, the learning dynamics are driven by the sum of per-task gradients: $G_{\text{total}}=\sum_{k=1}^{K}g_{k}$, where $g_{k}=\nabla_{\theta}\mathcal{L}^{k}$. The core problem, known as gradient conflict (Liu et al., 2021), arises when these per-task gradients are misaligned, a condition quantified by their cosine similarity: $\cos(g_{i},g_{j})<0$. A negative similarity indicates that an update improving performance on task $i$ actively degrades performance on task $j$. In a monolithic architecture where all core parameters are shared, frequent gradient conflicts lead to destructive interference (Sodhani et al., 2021). This dynamic forces the model into a suboptimal compromise, creating the representational bottlenecks and catastrophic performance drops characteristic of plasticity collapse (Dohare et al., 2024).
2.3 Architectural Paradigms for Mitigating Interference
To overcome the limitations of monolithic designs, architectural solutions have been proposed to enable parameter specialization and mitigate interference. Two prominent paradigms are particularly relevant. Sparse Mixture-of-Experts replaces the dense feed-forward network in a Transformer block with a set of $N$ parallel "expert" networks (Cai et al., 2025). For each input, a trainable gating function $G(x)$ sparsely selects a small subset of experts (e.g., top-k) to process the token: $\mathrm{MoE}(x)=\sum_{i=1}^{N}G_{i}(x)\cdot\mathrm{Expert}_{i}(x).$ This conditional computation creates specialized pathways within the model, allowing different inputs (e.g., from different tasks) to be processed by distinct subsets of parameters, thereby reducing interference and increasing model capacity without a proportional rise in computational cost. Recent designs also explore hybrid approaches, such as including a shared expert alongside specialized ones, to balance generalization and specialization (Dai et al., 2024). Low-Rank Adaptation offers a parameter-efficient method for adapting large pre-trained models (Hu et al., 2022). Instead of fine-tuning the entire weight matrix $W_{0}$, LoRA freezes $\theta_{B}$ and injects a trainable, low-rank "update" matrix, $\Delta\theta=BA$. The modified forward pass becomes $(\theta_{B}+\alpha BA)x$. By training only the small low-rank factors $(A,B)$, LoRA can efficiently learn task-specific modifications while preserving the general knowledge in the pre-trained weights. This approach has proven effective for creating extensive multitask capabilities on top of a single base model, where different LoRA modules can be composed to handle diverse tasks (Huang et al., 2023).
@@ 3 Method
We begin in Sec. 3.1 by quantitatively diagnosing plasticity collapse in multitask training. To address this, Sec. 3.2 presents a systematic design space exploration across five axes—which culminates in our proposed model, ScaleZero. Finally, to resolve the static allocation issue, we introduce Dynamic Parameter Scaling in Sect. 3.3, a LoRA-based strategy that adaptively allocates model capacity.
3.1 Diagnosing Plasticity Collapse in Multitask Planning